# AAAI Press Formatting Instructions for Authors Using — A Guide

### v2

@@ Appendix A Dataset
Dataset for Evaluation
Our datasets include real-world hospital datasets (BRSET (Nakayama et al. 2024) and VinDr‑PCXR (Pham, Tran, and Nguyen 2022)) alongside a variety of classification and visual question answering datasets. We focus on a low-resource data-efficient setting (training sets ranging from 546 to 7007 samples in classification, 1790-19,700 in VQA). Medical image annotation is a resource-intensive task, demanding expert annotators whose availability is limited and expensive. Clinical AI also often has poor generalization across heterogeneous medical centers and patient populations, often requiring that models be trained for the specific context in which they are deployed (Casey 2022). Such constraints often restrict researchers to datasets comprising only a few thousand samples for a given study. Retrieval augmentation is well-suited for this setting, as it is known to benefit low‑data regimes the most by leveraging external knowledge to compensate for sparse training samples. For the classification tasks, we used the following datasets:
BreastMNIST (nickname: Breast) (Al-Dhabyani et al. 2020) is a binary classification dataset of breast ultrasound. Following the official split, we use 546 for training, 78 for validation and 156 for testing.
RetinaMNIST (nickname: Retina) (Liu et al. 2022) is a multi-label classification dataset of retina Fundus Camera. Following the official split, we use 1,080 for training, 120 for validation and 400 for testing.
DermaMNIST (nickname: Derma) (Tschandl, Rosendahl, and Kittler 2018) is a multi-class classification dataset of common pigmented skin lesions. Following the official split, we use 7,007 for training, 1,003 for validation and 2,005 for testing. The dataset contains clinical images from 7 different diagnostic categories: actinic keratoses, basal cell carcinoma, benign keratosis, dermatofibroma, melanoma, melanocytic nevi, and vascular lesions.
VinDr-PCXR (Pham, Tran, and Nguyen 2022) is a large, open pediatric chest X-ray dataset collected in Vietnam (2020–2021) containing 9,125 posteroanterior scans from patients under 10 years old. It provides both lesion-level bounding-box annotations for 36 findings and image-level labels for multi-label of 15 diagnoses, curated by experienced radiologists. We follow the official split, which includes 7,728 training and 1,397 test images, with an additional internal split of the training data into 85% for training and 15% for validation.
BRSET (Nakayama et al. 2024) is a Brazilian multilabel ophthalmological dataset comprising retinal fundus images annotated with 14 distinct pathological findings. The dataset, composed of 16,266 images, presents a challenging real-world scenario for multilabel classification. There is no publicly available split; therefore, we apply an internal split for train, validation, and test sets (70% train, 10% validation, and 20% test).
Breast, Derma, and Retina taken from the large-scale MNIST-like collection of standardized biomedical images, including 12 datasets for 2D and 6 datasets for 3D. We especially used MedMNIST+ (Yang et al. 2023), which is a higher resolution extension of the original MedMNIST. We used the highest resolution of 224 × 224.
For medical question answering, we have three well-known datasets:
VQA-RAD (Lau et al. 2018) is a clinician-annotated visual question answering dataset focused on radiology images (e.g., X-ray, CT, MRI). It pairs each image with both open-ended and yes/no questions. Following the official split, we use 1,753 questions for training and 453 for testing. In addition, we further partition the training set into 85% for training and 15% for validation.
PathVQA (He et al. 2020) is a pathology-focused medical VQA dataset built from textbooks and the PEIR digital library, comprising 4,289 images and 32,632 question–answer pairs spanning both open-ended and yes/no questions. We follow the official split provided by the authors: 19,700 for training, 6,260 for validation, and 6,720 for testing.
SLAKE (Liu et al. 2021) is a bilingual medical VQA dataset. We use the English subset, SLAKE-English, which comprises 642 radiology images and 7.03k English QA pairs. We follow the official train/validation/test split of 4.92k/1.05k/1.06k QA pairs, and include all question types, covering both open-ended and closed-ended formats.
Dataset for Index
For construction of the Index we used three large datasets of (image, text) pairs: PMC-OA (Lin et al. 2023), ROCO (Rückert et al. 2024) and MIMIC-CXR (Johnson et al. 2019):
@@ Appendix B Implementation Details
Retrieval Augmentation
As described earlier, for each image–question pair, we retrieve $r$ image–text pairs, with $r=4$. We use the input image as the query and Jina-CLIP (Xiao, Mastrapas, and Wang 2024) as the retriever head in our multimodal retriever. The index is constructed with FAISS (Johnson, Douze, and Jégou 2019) over MIMIC-CXR, PMC-OA, and ROCO, which are fully described in Section A in Appendix A. To embed images, we use the Jina-CLIP visual head (the same model used for retrieval); to embed captions/reports, we use the Jina-CLIP text head. The index is stored in float16 due to storage constraints. We also experimented with BiomedCLIP (Zhang et al. 2023a).
Reader Fine-Tuning
We use Pixtral (12B) (Agrawal et al. 2024) and Qwen2-vl (7B) (Wang et al. 2024) as base models. We train the models with and without retrieval augmentation to assess their effect. The retrieval-augmented prompt is described in Appendix E. Note, we also tried Med-Flamingo (9B) (Moor et al. 2023) as a base model.
We fine-tune Pixtral and Qwen2-vl using LLaMA-Factory (Zheng et al. 2024) on a single NVIDIA L40 (48 GB) GPU for 10 epochs. We use a learning rate of $2\times 10^{-5}$ and apply Low-Rank Adaptation (LoRA) (Hu et al. 2021) for parameter-efficient fine-tuning. To all models (including baseline models), we conduct a grid search over batch size (2, 4, 6) and whether to freeze the vision head, selecting the best configuration by validation performance. For Med-Flamingo, we fine-tune using our codebase on a single NVIDIA L40 (48 GB) GPU for 10 epochs. Following the Med-Flamingo paper, the language model and image encoder are frozen, and only the Gated Cross-Attention layers and the Perceiver Resampler are optimized for stable and efficient learning. We use a learning rate of $2\times 10^{-5}$.
Multimodal Retriever Fine-Tuning
@@
$\mathrm{Prec}=\frac{|\hat{A}\cap A|}{|\hat{A}|},\quad\mathrm{Rec}=\frac{|\hat{A}\cap A|}{|A|},\quad\mathrm{F1}=\frac{2\,\mathrm{Prec}\cdot\mathrm{Rec}}{\mathrm{Prec}+\mathrm{Rec}}.$
@@
Backbone
BiomedCLIP
Variant
Breast Derma Retina Vindr BREST VQA-RAD SLAKE PathVQA Mean
@@
Qwen2-vl FT RAG .84 .81 .73 .38 .63 .45 .54 .09 .49 .27 .77 .89 .92 .72 .40
Jomed .87 .82 .75 .40 .65 .50 .58 .15 .50 .39 .79 .91 .92 .74 .45
Pixtral FT RAG .88 .84 .79 .57 .57 .46 .55 .09 .44 .31 .74 .87 .88 .72 .45
Jomed .91 .87 .80 .59 .59 .49 .56 .15 .50 .39 .77 .89 .89 .74 .50
@@ Appendix C Baseline Methods
An open multimodal generative pre-trained transformer for biomedicine that aligns diverse biological/biomedical modalities with natural language; we include it as a strong domain baseline (Luo et al. 2023).
@@
A biomedical adaptation of LLaVA trained via curriculum (biomedical figure–caption alignment followed by instruction tuning); we evaluate three commonly used releases as separate baselines (Li et al. 2023).
@@
Two implementations of Medical Visual Instruction Tuning from PMC-VQA (Zhang et al. 2023b). Both adopt a PMC-CLIP vision encoder (Lin et al. 2023). TE uses an encoder-style text pathway feeding a multimodal decoder, while TD concatenates text tokens with visual features to a decoder-only pathway; we report both as distinct baselines.
@@
MedDr is a generalist medical VLLM trained on large-scale instruction-style data curated from diagnosis-guided bootstrapping and medical image descriptions, covering multiple modalities and tasks (He et al. 2024a). MedDr+RAD augments MedDr at inference with Retrieval-Augmented Diagnosis (RAD): for a test image, similar cases are retrieved and summarized as contextual guidance in the prompt (He et al. 2024b). GSCo (Generalist–Specialist Collaboration) is a two-stage framework that (i) builds a generalist GFM (MedDr) and lightweight specialist models, and (ii) performs collaborative inference via Mixture-of-Expert Diagnosis (MoED; using specialist predictions as references) and RAD (using specialists to retrieve similar cases) (He et al. 2024b). GSCo evaluates across a large multi-dataset benchmark; we include GSCo as a strong InternVL-based baseline along with its MedDr and MedDr+RAD components (He et al. 2024b; Chen et al. 2024).
@@ Appendix D Additional Analysis
Performance with a State-of-the-Art LVLM Reranker
We evaluate whether a state-of-the-art large vision–language model (LVLM), used as an optional reranker, can improve performance on inconsistent-retrieval predictions. Specifically, we test the o3 reasoning model (OpenAI 2025) as a reranker: given a query image to classify and four retrieved image–caption pairs, o3 is asked to select the single pair that is most informative for predicting the correct label. We compare this reranking to four alternatives: (i) using the top-1 retrieved candidate (no reranking), (ii) selecting the model’s prediction with the highest confidence (maximum logit), (iii) using o3 with captions only (no images), and (iv) the Jomed aggregation strategy.
@@
Jomed variants with medical pre-trained backbones
Although our work focuses on general-purpose LVLMs, we also evaluated a medical-specific baseline. We conducted the evaluation on four benchmarks—Breast, Derma, Retina, and VQA-RAD (closed-question subset). For the medical pre-trained baseline, we used Med-Flamingo and paired it with a medically pre-trained retriever (BiomedCLIP), using only one retrieval head (image retrieval without text retrieval). The performance gap between Med-Flamingo and our reader approach was small—0.01 in accuracy and 0.03 in F1. We hypothesize that this narrow gap may be due to Med-Flamingo’s relatively older LLaMA backbone (Touvron et al. 2023). Results are presented in Table 10. We also tested our general-purpose backbone with a medically pre-trained retriever, which showed improvements over FT RAG, with a similar pattern to the general-purpose retriever. Full results are presented in Table 9.
Retrieval Robustness Analyses.
@@ Appendix E Prompt Design
Here, we detail the prompts used in our experiments and analysis. To explore the utility of o3 as re-rankers of the most predictive (image, caption) pair, as described in Section D in Appendix D, we provide both models with a detailed prompt for the task, shown in Prompt 1.
Additionally, we present the prompts used to train Jomed for each of the datasets: Prompts 2, 4, 3, 6, 5, and the prompt for visual question answering tasks Prompt 7.
1 Task:
2
3 You are an expert assistant selecting the most informative reference for medical image classification.
4
5 Description:
6 A patients [MRI/CT/X-ray] image needs to be classified into one of the following categories: [List of possible labels].
7 You are given four candidate reference pairs, each consisting of a caption and its associated image, drawn from PubMed literature.
8 Your goal is to determine which single candidate provides the most useful information to help an AI model correctly classify the provided image.
9
10 Input Information:
11 - Medical image: The patients MRI/CT/X-ray image is provided here.
12 - Classification task: Classify the image into one of the following categories: [List of possible labels].
13 - Candidates:
14 - Candidate 0:
15 - Caption: [Caption 0]
16 - Image: [Image 0]
17 - Candidate 1:
18 - Caption: [Caption 1]
19 - Image: [Image 1]
20 - Candidate 2:
21 - Caption: [Caption 2]
22 - Image: [Image 2]
23 - Candidate 3:
24 - Caption: [Caption 3]
25 - Image: [Image 3]
26
27 Instructions:
28 1. For each candidate, carefully assess how informative its caption and image are for the current classification task.
29 2. Specifically, evaluate whether the candidate describes imaging features, findings, or clinical context relevant to distinguishing among the listed categories.
30 3. Compare all candidates and select the one that gives the clearest, most discriminative information to aid the classification.
31 4. Output only the number of the selected candidate in the following format:
32 Caption Number: [your selected number]
Listing 1: Prompt used for o3
@@ 32 Caption Number: [your selected number]
Listing 2: Prompt used by Jomed for Breast classification
@@
Listing 3: Prompt used by Jomed for Retina classification
@@
Listing 4: Prompt used by Jomed for Derma classification
@@
Listing 5: Prompt used by Jomed for Vindr-PCXR classification
@@
Listing 6: Prompt used by Jomed for BRSET classification
@@
Listing 7: Prompt used by Jomed for visual question answering tasks

### v3

@@ Appendix B Implementation Details
B.1 Retrieval Augmentation
As described earlier, for each image–question pair, we retrieve $r$ image–text pairs, with $r=4$. We use the input image as the query and Jina-CLIP Xiao et al. (2024) as the retriever head in our multimodal retriever. The index is constructed with FAISS Johnson et al. (2019b) over MIMIC-CXR, PMC-OA, and ROCO, which are fully described in Section A.2 in Appendix A. To embed images, we use the Jina-CLIP visual head (the same model used for retrieval); to embed captions/reports, we use the Jina-CLIP text head. The index is stored in float16 due to storage constraints. We also experimented with BiomedCLIP Zhang et al. (2023a).
B.2 Reader Fine-Tuning
We use Pixtral (12B) Agrawal et al. (2024) and Qwen2-vl (7B) Wang et al. (2024) as base models. We train the models with and without retrieval augmentation to assess their effect. The retrieval-augmented prompt is described in Appendix E. Note, we also tried Med-Flamingo (9B) Moor et al. (2023) as a base model.
We fine-tune Pixtral and Qwen2-vl using LLaMA-Factory Zheng et al. (2024) on a single NVIDIA L40 (48 GB) GPU for 10 epochs. We use a learning rate of $2\times 10^{-5}$ and apply Low-Rank Adaptation (LoRA) Hu et al. (2021) for parameter-efficient fine-tuning. To all models (including baseline models), we conduct a grid search over batch size (2, 4, 6) and whether to freeze the vision head, selecting the best configuration by validation performance. For Med-Flamingo, we fine-tune using our codebase on a single NVIDIA L40 (48 GB) GPU for 10 epochs. Following the Med-Flamingo paper, the language model and image encoder are frozen, and only the Gated Cross-Attention layers and the Perceiver Resampler are optimized for stable and efficient learning. We use a learning rate of $2\times 10^{-5}$.
B.3 Multimodal Retriever Fine-Tuning
@@ B.4 Evaluation Metrics
$\mathrm{Prec}=\frac{|\hat{A}\cap A|}{|\hat{A}|},\quad\mathrm{Rec}=\frac{|\hat{A}\cap A|}{|A|},\quad$
$\mathrm{F1}=\frac{2\,\mathrm{Prec}\cdot\mathrm{Rec}}{\mathrm{Prec}+\mathrm{Rec}}.$
@@
Model
Breast Derma Retina Vindr BREST VQARAD SLAKE PathVQA Mean
@@
FT RAGQwen2-vl .84 .81 .73 .38 .63 .45 .54 .09 .49 .27 .77 .89 .92 .72 .40
JomedQwen2-vl .87 .82 .75 .40 .65 .50 .58 .15 .50 .39 .79 .91 .92 .74 .45
FT RAGPixtral .88 .84 .79 .57 .57 .46 .55 .09 .44 .31 .74 .87 .88 .72 .45
JomedPixtral .91 .87 .80 .59 .59 .49 .56 .15 .50 .39 .77 .89 .89 .74 .50
@@ Appendix C Baseline Methods
C.1 RAG baseline
We evaluate our method against several representative RAG-based and multimodal retrieval approaches to ensure a fair and comprehensive comparison.
RAD He et al. (2024a) is a retrieval-based method designed for classification tasks. During inference, the most similar image from the training set is retrieved, and its supervised label is directly used as the prediction. We integrate RAD with both of our backbone models (Qwen2-VL and Pixtral) under the same data splits and retrieval settings.
MMed-RAGXia et al. (2024a) represents a recent state-of-the-art multimodal RAG approach in which the retriever and LVLM are trained independently. We follow the official GitHub implementation, first training the CLIP module and then performing DPO training using two backbone models: LLaVA-Med and LLaVA.
Fusion-in-Decoder (FiD) Pipeline Izacard and Grave (2020) serves as a standard RAG-style baseline. In this setting, the retriever is frozen, and only the reader is fine-tuned while processing retrieved image–caption pairs as contextual input.
Perplexity Distillation Loss (PDist) Izacard et al. (2023). For each query and candidate document, we compute the reduction in perplexity (or equivalently, the increase in likelihood) of the correct output when the language model is conditioned on that document. From these likelihoods, we derive a “posterior” over documents (proportional to their contributions), and train the retriever to mimic that posterior via a KL divergence objective. In doing so, the retriever is guided to prioritize documents that most effectively support the language model’s generation, thereby aligning retrieval with generation quality.
C.2 Medical Pre - trained baseline models
@@
An open multimodal generative pre-trained transformer for biomedicine that aligns diverse biological/biomedical modalities with natural language; we include it as a strong domain baseline Luo et al. (2023).
@@
A biomedical adaptation of LLaVA trained via curriculum (biomedical figure–caption alignment followed by instruction tuning); we evaluate three commonly used releases as separate baselines Li et al. (2023).
@@
Two implementations of Medical Visual Instruction Tuning from PMC-VQA Zhang et al. (2023b). Both adopt a PMC-CLIP vision encoder Lin et al. (2023). TE uses an encoder-style text pathway feeding a multimodal decoder, while TD concatenates text tokens with visual features to a decoder-only pathway; we report both as distinct baselines.
@@
MedDr is a generalist medical VLLM trained on large-scale instruction-style data curated from diagnosis-guided bootstrapping and medical image descriptions, covering multiple modalities and tasks He et al. (2024a). MedDr+RAD augments MedDr at inference with Retrieval-Augmented Diagnosis (RAD): for a test image, similar cases are retrieved and summarized as contextual guidance in the prompt He et al. (2024b). GSCo (Generalist–Specialist Collaboration) is a two-stage framework that (i) builds a generalist GFM (MedDr) and lightweight specialist models, and (ii) performs collaborative inference via Mixture-of-Expert Diagnosis (MoED; using specialist predictions as references) and RAD (using specialists to retrieve similar cases) He et al. (2024b). GSCo evaluates across a large multi-dataset benchmark; we include GSCo as a strong InternVL-based baseline along with its MedDr and MedDr+RAD components He et al. (2024b); Chen et al. (2024).
@@ Appendix D Additional Analysis
D.1 Analysis of model components
We analyze the contribution of each part in the model in Table 7(a) and in Table 7(b). The result of the first stage of training - the reader retrieval augmentation fine-tuning is the FT RAG then we check the contribution of training only the text retriever head and then finally is the contribution of training both the text and image retriever head as Jomed. We also explore the utility of converting the open question into a closed one in our retrieval loss. If we don’t apply o3 model to convert the open question, we only trained on the closed one, which is referred to in Table 7(b) as Closed Question Only.
D.2 Performance with a State-of-the-Art LVLM Reranker
We evaluate whether a state-of-the-art large vision–language model (LVLM), used as an optional reranker, can improve performance on inconsistent-retrieval predictions. Specifically, we test the o3 reasoning model OpenAI (2025) as a reranker: given a query image to classify and four retrieved image–caption pairs, o3 is asked to select the single pair that is most informative for predicting the correct label. We compare this reranking to four alternatives: (i) using the top-1 retrieved candidate (no reranking), (ii) selecting the model’s prediction with the highest confidence (maximum logit), (iii) using o3 with captions only (no images), and (iv) the Jomed aggregation strategy.
@@ D.2 Performance with a State-of-the-Art LVLM Reranker
D.3 Jomed variants with medical pre-trained backbones
Although our work focuses on general-purpose LVLMs, we also evaluated a medical-specific baseline. We conducted the evaluation on four benchmarks—Breast, Derma, Retina, and VQA-RAD (closed-question subset). For the medical pre-trained baseline, we used Med-Flamingo and paired it with a medically pre-trained retriever (BiomedCLIP), using only one retrieval head (image retrieval without text retrieval). The performance gap between Med-Flamingo and our reader approach was small—0.01 in accuracy and 0.03 in F1. We hypothesize that this narrow gap may be due to Med-Flamingo’s relatively older LLaMA backbone Touvron et al. (2023). Results are presented in Table 11. We also tested our general-purpose backbone with a medically pre-trained retriever, which showed improvements over FT RAG, with a similar pattern to the general-purpose retriever. Full results are presented in Table 10.
D.4 Retrieval Robustness Analyses.
@@ Appendix E Prompt Design
We detail the prompts used in our experiments and analyses. To convert open questions into closed ones with the o3 model, we use the prompt in Prompt 1. To assess the utility of o3 as a re-ranker of the most predictive image–caption pair (Section D.2, Appendix D), we provide both LVLM backbones (Pixtral and Qwen2-VL) with the task-specific prompt in Prompt 2.
We also include the training prompts for Jomed across datasets—Prompts 3, 5, 4, 7, 6—and the VQA prompt (Prompt 8).
Convert the following open-ended question into a closed yes/no question based on the given answer.
The new question should be answerable with "{expected_answer}".
Original Question: {question}
Original Answer: {answer}
Expected New Answer: {expected_answer}
Please provide only the new closed question without any additional text or explanation.
Listing 1: Prompt used for o3 to convert open question into closed one
Task:
You are an expert assistant selecting the most informative reference for medical image classification.
Description:
A patients [MRI/CT/X-ray] image needs to be classified into one of the following categories: [List of possible labels].
You are given four candidate reference pairs, each consisting of a caption and its associated image, drawn from PubMed literature.
Your goal is to determine which single candidate provides the most useful information to help an AI model correctly classify the provided image.
Input Information:
- Medical image: The patients MRI/CT/X-ray image is provided here.
- Classification task: Classify the image into one of the following categories: [List of possible labels].
- Candidates:
- Candidate 0:
- Caption: [Caption 0]
- Image: [Image 0]
- Candidate 1:
- Caption: [Caption 1]
- Image: [Image 1]
- Candidate 2:
- Caption: [Caption 2]
- Image: [Image 2]
- Candidate 3:
- Caption: [Caption 3]
- Image: [Image 3]
Instructions:
1. For each candidate, carefully assess how informative its caption and image are for the current classification task.
2. Specifically, evaluate whether the candidate describes imaging features, findings, or clinical context relevant to distinguishing among the listed categories.
3. Compare all candidates and select the one that gives the clearest, most discriminative information to aid the classification.
4. Output only the number of the selected candidate in the following format:
Caption Number: [your selected number]
Listing 2: Prompt used for o3
@@ 4. Output only the number of the selected candidate in the following format:
Listing 3: Prompt used by Jomed for Breast classification
@@
Listing 4: Prompt used by Jomed for Retina classification
@@
Listing 5: Prompt used by Jomed for Derma classification
@@
Listing 6: Prompt used by Jomed for Vindr-PCXR classification
@@
Listing 7: Prompt used by Jomed for BRSET classification
@@
Listing 8: Prompt used by Jomed for visual question answering tasks